# On thermal transpiration and thermomolecular pressure difference

### v1

@@ 2.1. Estimates for $\nu$ and $K$
Recall that $L(f)=\nu(v)f-K(f)$ when the assumption (1.6) holds, and that $\nu(v)$ and $K(f)$ have explicit forms as follows
@@
$k_{2}(v,u)$ $=\frac{2B_{0}}{\pi^{3/2}}\frac{1}{|v-u|}e^{-\frac{1}{8}|v-u|^{2}-|V_{1}(v,u)|^{2}}\times\int_{W_{v-u}}e^{-|w+V_{2}(v,u)|^{2}}\left(|v-u|^{2}+|w|^{2}\right)^{\frac{\gamma}{2}}dw,$
@@
Suppose that (1.6) holds and $3<\gamma\leq 1$. For the collision frequency $\nu(v)$, there exist positive constant $\nu_{0}$ and $\nu_{1}$ such that
@@
Let $3<\gamma\leq 1$ and let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$. Suppose (1.6) holds. Then, we have
@@
Let $3<\gamma\leq 1$ and $0<\delta<1$. Suppose that (1.6) holds. Then, we have
$|k(v,u)|\leq C_{\gamma,\delta}\frac{w_{\gamma}(|v-u|)}{(1+|v|+|u|)^{1-\gamma}}E_{\delta}(v,u)$ (2.1)
@@
$w_{\gamma}(|v-u|)\coloneq\left\{\begin{aligned} &\frac{1}{|v-u|},&&-1<\gamma\leq 1,\\ &\frac{|\ln|v-u||+1}{|v-u|},&&\gamma=-1,\\ &\frac{1}{|v-u|^{|\gamma|}},&&-3<\gamma<-1.\end{aligned}\right$. (2.2)
@@
In the chapter 4, we estimate $L^{\infty}_{\alpha,0}$ norm of the solution of Boltzmann equation. Hence, we need to discuss
@@
Plugging these identities into that estimate 2.1 with $a=\alpha$ and $\delta=1/2-\alpha$, we obtain
$|k_{\alpha}(v,u)|\leq C_{\gamma,\alpha}\frac{w(|v-u|)}{(1+|v|+|u|)^{1-\gamma}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}$ (2.3)
@@
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}\leq\frac{C_{\alpha,\mu}}{1+|v|}$ (2.4)
@@
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}dv$ $=\int_{0}^{\infty}\int_{0}^{2\pi}\int_{0}^{\pi}\frac{1}{r^{\mu-2}}e^{-\alpha_{1}r^{2}}e^{-\alpha_{3}\left(|v|\cos\theta-\alpha_{2}r\right)^{2}}\sin\theta d\theta d\phi dr$
@@
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}\leq C.$
Here, the constant $C$ is independent to $v$. combining these results, the proof is complete. ∎
@@
[24] Let $0\leq\alpha<1/4,\beta\in\mathbb{R}^{3},-3<\gamma\leq 1$ and $1\leq q<\min\{3,\frac{3}{|\gamma|}\}$. Suppose that (1.6) holds. Then we have
$\displaystyle\int_{\mathbb{R}^{3}}|k_{\alpha}(v,u)|^{q}(1+|u|)^{\beta}du\leq C_{\alpha,\gamma}(1+|v|)^{\beta+q(\gamma-1)-1}$ (2.5)
@@
Notice that $(1+|u|)^{\beta}\leq C_{\beta}\left((1+|v|)^{\beta}+|v-u|^{\beta}\right)$ for all $v,u\in\mathbb{R}^{3}$. By 2.3, we have
@@
$w_{-1}(|v-u|)e^{-\alpha_{1}|v-u|^{2}}\leq C_{\epsilon}|v-u|^{-(1+\epsilon)}$ (2.6)
for some small positive constant $\epsilon$. Thus, we can treat the case $\gamma=-1$ in the same way as for $3<\gamma<-1$. In any case, by (2.4) we have
$\displaystyle\int_{\mathbb{R}^{3}}|k_{\alpha}(v,u)|^{q}(1+|u|)^{\beta}du$ $\leq C_{\alpha,\gamma}(1+|v|)^{q(\gamma-1)-1}+C_{\alpha,\gamma}(1+|v|)^{\beta+q(\gamma-1)-1}$
@@ 2.2. Proposition of $L$
$\langle L(f),f\rangle_{v}\geq C_{L}\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}_{0,\gamma/2}(dv)}.$ (2.7)
Here $L^{2}_{0,\gamma/2}(dv)$ denotes integration only with respect to the variable $v$. The proof can be found in [17], [20] and [22]. The most important estimate concerning $P_{L}$ and $(I-P_{L})$ is the following.
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}f+\frac{1}{\kappa}L(f)&=\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f&=P(f)+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2}\end{aligned}\right$. (2.8)
@@
$\lVert P_{L}(f)\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\leq C\begin{pmatrix}[l]&\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}+\lVert(I-P)f\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}\\ &+\lVert f\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}\end{pmatrix}.$ (2.9)
@@
Recall $P_{L}(f)=\sum_{i=0}^{4}\langle f,\Phi_{i}\rangle_{v}\Phi_{i}$. Thus, it suffices to compute $\langle f,\Phi_{i}\rangle$, for $i=0,\cdots,4$. The estimate of $\langle f,\Phi_{i}\rangle$ are the same as those given in lemma 3.3 in [14] for $i=1,\cdots,4$. The standard elliptic estimates appearing in [14] can also be found in the chapter 2 of [18] or chapter 6 of [15]. For $i=0$, we choose a test function
$\psi=\sum_{j=1}^{3}(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v_{j}\partial_{j}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0},$ (2.10)
@@
$\left\{\begin{aligned} -\Delta\phi_{\langle f,\Phi_{0}\rangle_{v}}&=\langle f,\Phi_{0}\rangle_{v},&&\text{in }\Omega,\\ \hat{\gamma}(x)\phi_{\langle f,\Phi_{0}\rangle_{v}}+\partial_{n}\phi_{\langle f,\Phi_{0}\rangle_{v}}&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (2.11)
Here $\hat{\gamma}(x)$ is a non-negative $C^{2}$ function and $\hat{\gamma}(x)=0$ on $\partial\Omega_{2}$ ( but $\gamma>0$ on a positive measure subset of $\partial\Omega$ ). Notice that we have following estimate
$\lVert\phi_{\langle f,\Phi_{0}\rangle_{v}}\rVert_{H^{2}(\Omega)}\leq C\lVert\langle f,\Phi_{0}\rangle_{v}\rVert_{L^{2}(\Omega)}.$ (2.12)
This estimate can be found in [16] and [34]. We choose $\beta_{\langle f,\Phi_{0}\rangle_{v}}$ such that
$\int_{\mathbb{R}^{3}}\left(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}}\right)\left(\frac{|v|^{2}}{2}-\frac{3}{2}\right)(v_{i})^{2}\,\mu(v)\,dv=0.$
Then, following the same discussion as in [14], we obtain
$\lVert\langle f,\Phi_{0}\rangle_{v}\rVert^{2}_{L^{2}(\Omega)}\leq C\begin{pmatrix}[l]&\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}(\Omega\times\mathbb{R}^{3})}+\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,\gamma/2}(\Omega\times\mathbb{R}^{3})}+\lVert(I-P)(f)\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}\\ &+\lVert f\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}\end{pmatrix}.$ (2.13)
Combining all these estimates, the proof is completed. ∎

### v2

@@ 2.1. Estimates for $\nu$ and $K$
Recall that $L(f)=\nu(v)f-K(f)$ when the assumption (8) holds, and that $\nu(v)$ and $K(f)$ have explicit forms as follows
@@
$k_{2}(v,u)$ $=\frac{2\mathfrak{B}}{\pi^{3/2}}\frac{1}{|v-u|}e^{-\frac{1}{8}|v-u|^{2}-|V_{1}(v,u)|^{2}}\times\int_{W_{v-u}}e^{-|w+V_{2}(v,u)|^{2}}\left(|v-u|^{2}+|w|^{2}\right)^{\frac{\gamma}{2}}dw,$
@@
Suppose that (8) holds and $3<\gamma\leq 1$. For the collision frequency $\nu(v)$, there exist positive constants $\nu_{0}$ and $\nu_{1}$ such that
@@
Let $3<\gamma\leq 1$ and let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$. Suppose (8) holds. Then, we have
@@
Let $3<\gamma\leq 1$ and $0<\delta<1$. Suppose that (8) holds. Then, we have
$|k(v,u)|\leq C_{\gamma,\delta}\frac{w_{\gamma}(|v-u|)}{(1+|v|+|u|)^{1-\gamma}}E_{\delta}(v,u)$ (27)
@@
$w_{\gamma}(|v-u|)\coloneq\left\{\begin{aligned} &\frac{1}{|v-u|},&&-1<\gamma\leq 1,\\ &\frac{|\ln|v-u||+1}{|v-u|},&&\gamma=-1,\\ &\frac{1}{|v-u|^{|\gamma|}},&&-3<\gamma<-1.\end{aligned}\right$. (28)
@@
In the section 4, we estimate $L^{\infty}_{\alpha,0}$ norm of the solution of Boltzmann equation. Hence, we need to discuss
@@
Plugging these identities into that estimate (27) with $a=\alpha$ and $\delta=1/2-\alpha$, we obtain
$|k_{\alpha}(v,u)|\leq C_{\gamma,\alpha}\frac{w(|v-u|)}{(1+|v|+|u|)^{1-\gamma}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}$ (29)
@@
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}\leq\frac{C_{\alpha,\mu}}{1+|v|}$ (30)
@@
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}$ $=\int_{0}^{\infty}\int_{0}^{2\pi}\int_{0}^{\pi}\frac{1}{r^{\mu-2}}e^{-\alpha_{1}r^{2}}e^{-\alpha_{3}\left(|v|\cos\theta-\alpha_{2}r\right)^{2}}\sin\theta d\theta d\phi dr$
@@
$\int_{\mathbb{R}^{3}}\frac{1}{|v-u|^{\mu}}e^{-\alpha_{1}|v-u|^{2}}e^{-\alpha_{3}\left(\frac{(v-u)\cdot v}{|v-u|}-\alpha_{2}|v-u|\right)^{2}}\leq C_{\alpha}.$
Here, the constant $C_{\alpha}$ is independent to $v$. Combining these results, the proof is complete. ∎
@@
[24] Let $0\leq\alpha<1/4,\beta\in\mathbb{R}^{3},-3<\gamma\leq 1$ and $1\leq q<\min\{3,\frac{3}{|\gamma|}\}$. Suppose that (8) holds. Then we have
$\displaystyle\int_{\mathbb{R}^{3}}|k_{\alpha}(v,u)|^{q}(1+|u|)^{\beta}du\leq C_{\alpha,\gamma}(1+|v|)^{\beta+q(\gamma-1)-1}$ (31)
@@
Notice that $(1+|u|)^{\beta}\leq C_{\beta}\left((1+|v|)^{\beta}+|v-u|^{\beta}\right)$ for all $v,u\in\mathbb{R}^{3}$. By (29), we have
@@
$w_{-1}(|v-u|)e^{-\alpha_{1}|v-u|^{2}}\leq C_{\epsilon}|v-u|^{-(1+\epsilon)}$ (32)
for some small positive constant $\epsilon$. Thus, we can treat the case $\gamma=-1$ in the same way as for $3<\gamma<-1$. In any case, by proposition 2.4 we have
$\int_{\mathbb{R}^{3}}|k_{\alpha}(v,u)|^{q}(1+|u|)^{\beta}du$ $\leq C_{\alpha,\gamma}(1+|v|)^{q(\gamma-1)-1}+C_{\alpha,\gamma}(1+|v|)^{\beta+q(\gamma-1)-1}$
@@ 2.2. Proposition of $L$
$\langle L(f),f\rangle_{v}\geq C_{L}\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}_{0,\gamma/2}(dv)}.$ (33)
Here, $L^{2}_{0,\gamma/2}(dv)$ denotes integration only with respect to the variable $v$. The proof can be found in [17], [20] and [22]. We present a key estimate concerning $P_{L}$ and $(I-P_{L})$.
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}f+\frac{1}{\kappa}L(f)&=\phi_{1},&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=f_{in},&&\quad\text{on}\quad\partial\Omega^{-}_{1},\\ f&=P(f)+\phi_{2}(x,v),&&\quad\text{on}\quad\partial\Omega^{-}_{2}\end{aligned}\right$. (34)
@@
$\lVert P_{L}(f)\rVert^{2}_{L^{2}_{0,\gamma/2}}$ $\leq C\begin{pmatrix}[l]&\left[1+\frac{1}{\kappa}\right]\lVert(I-P_{L})(f)\rVert^{2}_{L^{2}_{0,\gamma/2}}+\lVert\phi_{1}\rVert^{2}_{L^{2}_{0,-\gamma/2}}+\lVert(I-P)f\rVert^{2}_{L^{2}(\partial\Omega^{+}_{2})}\\ &+\lVert f\rVert^{2}_{L^{2}(\partial\Omega^{+}_{1})}+\lVert f_{in}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{1})}+\lVert\phi_{2}\rVert^{2}_{L^{2}(\partial\Omega^{-}_{2})}\end{pmatrix}.$ (35)
@@
Recall $P_{L}(f)=\sum_{i=0}^{4}\langle f,\Phi_{i}\rangle_{v}\Phi_{i}$. Thus, it suffices to compute $\langle f,\Phi_{i}\rangle_{v}$, for $i=0,\cdots,4$. The estimate of $\langle f,\Phi_{i}\rangle_{v}$ are the same as those given in lemma 3.3 in [14] for $i=1,\cdots,4$. The standard elliptic estimates appearing in [14] can also be found in the section 2 of [18] or section 6 of [15]. For $i=0$, due to the difference between the mixed and diffuse boundary conditions we need to find a different test function to estimate $\langle f,\Phi_{0}\rangle$. The Green’s identity provides the following weak version of (34):
$\int_{\partial\Omega\times\mathbb{R}^{3}}\psi f\,dSdv-\iint_{\Omega\times\mathbb{R}^{3}}v\cdot\nabla_{x}\psi\,fdxdv=-\iint_{\Omega\times\mathbb{R}^{3}}\psi\,\frac{1}{\kappa}L(\mathrm{I}-\mathbf{P})fdxdv+\iint_{\Omega\times\mathbb{R}^{3}}\psi\,\phi_{1}dxdv.$ (36)
We choose the test function
$\psi=(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v\cdot\nabla_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0},$ (37)
@@
$\left\{\begin{aligned} -\Delta\phi_{\langle f,\Phi_{0}\rangle_{v}}&=\langle f,\Phi_{0}\rangle_{v},&&\text{in }\Omega,\\ \chi(x)\phi_{\langle f,\Phi_{0}\rangle_{v}}+\partial_{n}\phi_{\langle f,\Phi_{0}\rangle_{v}}&=0,&&\text{on }\partial\Omega.\end{aligned}\right$. (38)
Here $\chi(x)$ is a non-negative $C^{2}$ function. We require that $\chi(x)=0$ on $\partial\Omega_{2}$ but $\chi>0$ on a positive measure subset of $\partial\Omega$. Notice that we have following estimate
$\lVert\phi_{\langle f,\Phi_{0}\rangle_{v}}\rVert_{H^{2}(\Omega)}\leq C\lVert\langle f,\Phi_{0}\rangle_{v}\rVert_{L^{2}(\Omega)}.$ (39)
This estimate can be found in [16] and [34]. We then bound the right hand side of (36) by
We choose $\beta_{\langle f,\Phi_{0}\rangle_{v}}$ such that
$\int_{\mathbb{R}^{3}}\left(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}}\right)\left(\frac{|v|^{2}}{2}-\frac{3}{2}\right)(v_{i})^{2}\,\mu(v)\,dv=0.$ (40)
Moreover, we decompose that
$\begin{aligned} f&=\mathbf{1}_{\partial\Omega^{+}_{1}}f+\mathbf{1}_{\partial\Omega^{-}_{1}}f_{in}+\mathbf{1}_{\partial\Omega_{2}}Pf+\mathbf{1}_{\partial\Omega^{+}_{2}}(I-P)f+\mathbf{1}_{\partial\Omega^{-}_{2}}\phi_{2},&&\qquad\text{on }\partial\Omega\times\mathbb{R}^{3},\\ f&=\sum_{i=0}^{4}\langle f,\Phi_{i}\rangle_{v}\Phi_{i}+(\mathrm{I}-\mathbf{P}_{L})f,&&\qquad\text{on }\Omega\times\mathbb{R}^{3}.\end{aligned}$ (41)
By (41), since the contribution of $\langle f,\Phi_{4}\rangle_{v}$ vanishes in (36) due to our choice of $\beta_{\langle f,\Phi_{0}\rangle_{v}}$ and the contributions of $\langle f,\Phi_{1}\rangle_{v}$, $\langle f,\Phi_{2}\rangle_{v}$, and $\langle f,\Phi_{3}\rangle_{v}$ vanish in (36) due to the oddness, the left-hand side of (36) takes the form of
$\int_{\partial\Omega\times\mathbb{R}^{3}}n\cdot v\left(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}}\right)v\cdot\nabla_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0}\bigl[\mathbf{1}_{\partial\Omega^{+}_{1}}f+\mathbf{1}_{\partial\Omega^{-}_{1}}f_{in}+\mathbf{1}_{\partial\Omega_{2}}Pf+\mathbf{1}_{\partial\Omega^{+}_{2}}(I-P)f+\mathbf{1}_{\partial\Omega^{-}_{2}}\phi_{2}\bigr]$ (42)
$-\sum_{i,k=1}^{d}\iint_{\Omega\times\mathbb{R}^{3}}(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v_{i}v_{k}\,\partial_{ik}\phi_{\langle f,\Phi_{0}\rangle_{v}}\,\langle f,\Phi_{0}\rangle_{v}\Phi^{2}_{0}dxdv$ (43)
$-\sum_{i,k=1}^{d}\iint_{\Omega\times\mathbb{R}^{3}}(|v|^{2}-\beta_{\langle f,\Phi_{0}\rangle_{v}})v_{i}v_{k}\,\partial_{ik}\phi_{\langle f,\Phi_{0}\rangle_{v}}\Phi_{0}\,(\mathrm{I}-\mathbf{P}_{L})f.$ (44)
$v=(v\cdot n)n+v-(v\cdot n)n=v_{n}n+v_{\perp}.$
The contribution of $Pf=z(x)m^{\frac{1}{2}}=z(x)\Phi_{0}$ is
Notice that the boundary condition (38) reduces to the Neumann boundary condition on $\partial\Omega_{2}$. Therefore, the first term vanish. While, the second term also vanishes due to the oddness of $v_{n}v_{\perp}$. Therefore, (42) and (44) are bounded by
and
$\|\phi_{\langle f,\Phi_{0}\rangle_{v}}\|_{H^{2}(dx)}\left(\lVert(I-P)f\rVert_{L^{2}(\partial\Omega^{+}_{2})}+\lVert f\rVert_{L^{2}(\partial\Omega^{+}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{2})}+\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}\right).$
For (43), if $k\neq i$, it vanishes due to the oddness. Hence, the only contribution comes from $k=i$.
Using $\Delta_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}={\langle f,\Phi_{0}\rangle_{v}}$ and $\lVert\phi_{\langle f,\Phi_{0}\rangle_{v}}\rVert_{H^{2}(dx)}\leq C\lVert\langle f,\Phi_{0}\rangle_{v}\rVert_{L^{2}(dx)}$, we obtain
We therefore deduce the proposition. ∎